# Humanoid Locomotion and Manipulation: Current Progress and Challenges in Control, Planning, and Learning thanks: *co-corresponding authors

### v1

@@ VII-B Skill Learning: Imitation from Robot Experience
Imitation Learning (IL) is an umbrella term that represents a class of algorithms, including supervised, unsupervised, and reinforcement learning, that train policies from expert demonstrations. IL is particularly effective for complex tasks that are difficult to specify explicitly. Three essential steps exist in IL [299]. The first step is to capture the expert demonstration. The next step involves retargeting, where these demonstrations are mapped to the robot motions. The final step is policy training using the retargeted data. If the captured motion comes from the same robot, such as from teleoperation, the retargeting step is unnecessary.
We discuss four possible sources of demonstrations for humanoid robots: (i) policy execution, (ii) teleoperation, (iii) motion capture, and (iv) human videos, as illustrated in Fig. 10 and Table III. We group these data sources into two categories: The first is robot experience data, representing those directly obtained from the robots through policy execution or teleoperation; The second is human data, which includes human motion captures and videos of human activities obtained from the Internet. Robot experiences exhibit smaller morphological discrepancies and are directly applicable to policy learning but are typically scarce. Conversely, human data are more abundant, but they present significant morphological differences to humanoid robots. This section discusses robot experience data, and the next subsection discusses the human data.
@@ VII-B1 Obtaining Robot Experience Data
A reliable way to obtain robot experience data is to execute existing expert policy, either model-based or learning-based. However, collecting data on a physical robot requires a laborious setup of the environment and raises significant safety concerns. Therefore, conducting these executions in simulation is a more efficient approach, although the fidelity of the simulator will inevitably cause a sim-to-real gap.
@@
Teleoperation is one of the most common ways to directly capture robot data commanded by human experts. A main advantage of teleoperation is its ability to provide smooth, natural, and precise trajectories for a wide range of tasks. Fig. 11 outlines the control flow of teleoperation data used as a source of policy training. The first step of this process is generating the demonstration through teleoperation, represented by the top path of the control flow (dashed lines). Motion retargeting maps human measurements from the teleoperation device to the desired trajectories in the robot domain. Robot execution data collected from teleoperation can be used to train autonomous policies that directly command the robot’s motion without human intervention (solid lines). In this case, the retargeting process is not needed.
However, teleoperation for the full-body of a humanoid has a number of limitations. Firstly, a majority of teleoperation systems capture only manipulation skills [8]. Generally, Virtual Reality (VR)-based teleoperation schemes cannot sense the operator’s gait and are restricted to simply commanding walking speeds and directions via joysticks [246, 300, 301]. Full-body sensing, including human gaits, often requires additional equipment such as IMU suits [244, 302] or exoskeletons [249]. These devices may be prohibitively expensive, bulky, complicated to maintain, and lack transparency and user-friendliness. Secondly, while teleoperation can generate versatile training data for a wide range of tasks, the utility of this data may be limited if the robot’s kinematics do not enable seamless retargeting of reference human motion. Additionally, retargeting dynamic tasks such as walking or pushing an object is sensitive to the dynamic model of the human demonstrator [244] and requires meeting high synchronicity [302] and rich sensory feedback.
Although both teleoperation data and policy execution data are recorded from robots, they exhibit distinct characteristics. During teleoperation, human instructors tend to provide diverse demonstrations even for the same task. Therefore, teleoperation data are often multimodal, i.e., given a specific task, there exists a distribution of plausible ways to accomplish the task. In contrast, data from executing a single policy are often unimodal, i.e., given an input, the output is often fixed. Different policy learning approaches have been proposed to address these multimodal and unimodal data features.
@@ VII-B2 Approaches to Learning from Robot Experience Data
From unimodal policy execution data, which contains paired observations and actions, IL approaches are often used for policy distillation. Behavior Cloning (BC) casts IL as a supervised learning problem, which remains one of the most straightforward approaches for robot skill learning [303, 304]. Another IL technique is Inverse Reinforcement Learning (IRL), which reconstructs rewards from the data in addition to training an RL policy. The IRL study in [290] infers a generalizable reward of the expert demonstration for bipedal locomotion and then uses this reward to train an RL policy in unseen terrains.
To capture the data multimodality and produce diverse future actions, Action Chunking Transformer (ACT) [272, 291] is adopted to handle distribution shifts due to the compounding error inherent in naive BC. Recently, diffusion policy [276], a BC method, shows the ability to acquire multimodal locomotion skills by learning from a large dataset collected from multiple expert policies. However, obtaining these skills requires large-scale versatile data, which motivates the scaling of data collection via teleoperation in industrial companies, e.g., Tesla and Toyota Research [8].
Conclusion: Although collecting high-quality data demands considerable effort and resources, IL from robot experience remains a reliable method for attaining skills with expert-level performance. Industrial companies and research labs are increasingly focusing on scaling data collection to develop a broader range of diverse policies through IL. Especially, teleoperation is one of the most popular ways to collect humanoid robot experiences nowadays. For further reading on collecting robot experience data, we recommend the survey on humanoid robot teleoperation [305]. We also find the survey on imitation learning of humanoid bipedal locomotion [299] a decent summary.
@@ VII-C Skill Learning: Imitation from Human Data
While robot experiences can serve as a reliable data source, collecting loco-manipulation data directly from robots remains a formidable challenge. Collecting teleoperation data, even though it is one of the most commonly used approaches, is costly and time-consuming to scale. Besides, gathering robot data by deploying existing model-based methods or trained policies presents additional difficulties. These methods raise safety concerns on hardware or suffer from significant sim-to-real gaps. Furthermore, model-based methods based on human knowledge (e.g., dynamics models, heuristic trajectories) tend to generate consistent but similar behaviors, leading to limited data diversity.
Learning from a large, diverse corpus of human data mitigates these challenges. Humans use loco-manipulation skills in their daily lives almost reflexively. Therefore, recording human data is more accessible and scalable. Training policies to imitate human data can greatly simplify the synthesis of loco-manipulation behaviors. Recent research efforts in 3D human motion data archival have surged in the vision and computer graphics communities. As shown in Fig. 12, there are currently two primary approaches to acquire 3D human motion data: (i) recording directly from motion capture systems and (ii) reconstructing from 2D videos.
@@ VII-C1 Obtaining Human Data
Various tracking systems are used to obtain human motion data. As shown in Fig. 13[306, 179] captures humans interacting with various objects while moving around. The following datasets, CMU [307], SFU [308], LAFAAN [309], and AMASS [310], are commonly used because they provide a wide variety of human motions. However, motion capture data often require heavily instrumented environments and actors, making them expensive to scale. Besides, indoor lab settings provide little exposure to outdoor activities.
Alternatively, videos and images obtained from the Internet offer a rich and diverse source of human motion data, including athletic performances, artistic dances, or daily chores. However, compared to the aforementioned data sources, motion extracted from internet data is usually of lower quality, containing noise, jittery, and non-physical artifacts due to occlusion and motion blurs. Therefore, the reconstruction of accurate 3D human poses from 2D data remains an active research topic in the computer vision community.
@@
Animation is also a widely explored approach to obtaining human motion data. Although animation is effective in designing expressive motions for virtual human characters, the process often requires the use of sophisticated animation tools by professional animators [311], which makes it less scalable than motion capture or Internet videos. To address this limitation, researchers have been pursuing motion generation, leveraging human data to generate diverse and realistic motions; see a comprehensive survey in [312]. Our focus here is on training IL policies to generate physically plausible actions that achieve the motions demonstrated in human data. The training pipeline is shown in Fig. 12.
@@ VII-C2 Challenges in Learning from Human Data
Using offline human motion data (from any source) to train humanoids inevitably creates an embodiment gap both in observation and action spaces due to different body proportions, joint configurations, and mass distributions between most humanoids and humans. Closing this embodiment gap requires retargeting, which involves mapping the motion collected from a source skeletal model to a target robot model. Previous work in computer graphics and robotics has explored various retargeting strategies, such as joint correspondence [244, 272, 313], task-space correspondence [291], contact points [221], fingertips [314], gait synchronization [249, 302], and motion feasibility filter [271, 156]. Developing systematic solutions for retargeting the entire human body, including dexterous hands, remains a critical topic for advancing humanoids.
Another correspondence problem is that these human data are only proprioception-based, which lacks sensory input and action output. Notably, these human data lack tactile or force measurement from interactions, which limits the capability of learning for loco-manipulation with rich physical interaction. To solve this issue, IL trains control policies that track references within a physical simulator. Specifically, IL control policies accept state-only references instead of state-action pairs and output control signals. The physical simulator plays a key role in providing sensory input and validating the physical feasibility of the policy action.
@@ VII-C3 Approaches to Learning from Human Data
Examples of learned human-like motions include walking using Generative Adversarial Imitation Learning (GAIL) [297, 293]. Extending the GAIL framework, Adversarial Motion Prior (AMP) [315] is applied to locomotion tasks in [294] and [295]. Recently, RL-based motion imitation using motion capture data has shown successful transfer to humanoid robots [313, 271, 272, 316]. However, most of these works achieve only relatively conservative loco-manipulation behaviors. Highly agile behaviors, as shown in DeepMimic [317], still exist only in simulators, and similar capabilities have yet to be replicated with real robots.
Human motion imitation can also achieve robust policies capable of rich interactions with objects in unstructured environments. For example, [296, 318] demonstrates the learning of full-body gymnastic skills for humanoids in physics simulation, utilizing video reconstruction data and RL-based motion imitation. [319, 320] learns loco-manipulation policies enabling simulated humanoid characters to carry boxes. Furthermore, [292] mimics motion capture data to achieve playing basketball and grasping objects. However, many of these approaches still rely on privileged information, such as the ground truth of object and ego poses, from the simulator, which limits their extension to real-world hardware.
Conclusion: Transferring skills from humans, especially through Internet-scale datasets, unlocks a broad range of loco-manipulation capabilities for humanoid robots. Although learning from human data holds great promise, a large body of what has been achieved in simulation has yet to be realized in real-world robots. Developing affordable and capable humanoid robots and high-fidelity simulators of real-world loco-manipulation scenarios can accelerate progress in this research direction. As we witness an increasing accessibility of humanoid robots in the market, we foresee that imitation from human data will enable humanoid robots to acquire a vast and diverse skill set. This skill set is fundamental to building humanoid foundation models discussed in Sec. VIII.
@@ VII-D1 Combining Pure RL and IL
The combination of IL and pure RL without demonstration data has led to effective sim-to-real transfer. A two-stage teacher-student paradigm is widely adopted [37, 39, 321]. In these works, a teacher policy is first trained from simulated privileged observation using pure RL. Then, a student policy clones the behavior of the teacher, achieving a similar performance using only partial observations. The trained student policy is readily deployable on hardware with onboard observations. Another two-stage paradigm [322] reverses the order of the two policies. It uses IL first to pre-train an imitation policy from expert data. Then, an RL policy fine-tunes the imitation policy for further improvement. This RL allows achieving performance beyond the IL expert and adapting to different environments or tasks [323].
@@ VII-D2 Learning to Track Trajectory Reference
Combining model-based methods with learning-based methods is extensively explored. Numerous works on quadruped robots [324, 325] use MPC to generate reference motions and use them as imitation rewards for RL-based motion imitation. However, calculating MPC online can prolong the training time and MPC may occasionally fail to find a feasible solution. Online RL that tracks offline-generated trajectories avoids this problem and has achieved effective RL policies for versatile bipedal locomotion [326, 35, 327]. Supervised learning is also used to mimic TO-generated offline motions. For example, [328] allows humanoid robots to achieve brachiation on monkey bars. In addition to enhancing learning methods with model-informed trajectories, a learned policy can also suggest good warm starts or hyperparameters for model-based methods, thus significantly reducing the iterations required for convergence to an optimal trajectory.
@@ VII-D3 Learning to Augment Trajectory Reference
Rather than tracking the reference trajectory, augmenting the reference with a residual is another popular approach. Early work on dynamic movement primitives [329] modifies the reference trajectory by learning a task-specific force output, which achieves a humanoid racket swing. A milestone in trajectory augmentation is the demonstration of bipedal locomotion in [330]. Since then, many works on trajectory augmentation have achieved dynamic locomotion [331, 332]. Recently, loco-manipulation skills are also achieved through augmenting TO-based reference trajectory [333]. Beyond joint-level reference, a policy can also augment the task-space reference. For example, [334] learn a locomotion policy to adapt the foot placement reference derived from inverted pendulum models.
Both imitation and augmentation of the reference trajectory share benefits and drawbacks. As a benefit, using references expedites the learning process and provides an effective way of learning complex skills. However, both methods rely on predefined trajectories and, therefore, have limited potential to learn emergent, diverse behaviors.
Conclusion: Learning-based methods, especially pure RL without demonstration data, present robust and emergent behaviors, while IL enables effective learning of complex behaviors. Model-based methods take advantage of human knowledge and the power of numerical optimization to provide references for efficient learning. Overall, combining model-based and learning-based approaches has achieved efficient, versatile, and high-performing humanoid tasks that outperform single methods alone. Many existing works have shown successful humanoid hardware deployments using hybrid methods.

### v2

@@ VII-B Skill Learning: Imitation from Robot Experience
Imitation Learning (IL) is an umbrella term that represents a class of algorithms, including supervised, unsupervised, and reinforcement learning, that train policies from expert demonstrations. IL is particularly effective for complex tasks that are difficult to specify explicitly. Three essential steps exist in IL [18]. The first step is to capture the expert demonstration. The next step involves retargeting, where these demonstrations are mapped to the robot motions. If the captured motion comes from the same robot, such as from teleoperation, the retargeting step is unnecessary. The final step is policy training using the retargeted data.
We discuss four possible sources of demonstrations for humanoid robots: (i) policy execution, (ii) teleoperation, (iii) motion capture, and (iv) human videos, as illustrated in Fig. 10 and Table V. We group these data sources into two categories: the robot experience data, which represents those directly obtained from the robots through policy execution or teleoperation, and the human data, which includes human motion captures and videos of human activities obtained from the Internet. Robot experiences exhibit smaller morphological discrepancies and are directly applicable to policy learning but are typically scarce. Conversely, human data are more abundant but present significant morphological differences to humanoid robots.
@@ VII-B1 Obtaining Robot Experience Data
A reliable way to obtain robot experience data is to execute existing policies, either model-based or learning-based. However, collecting data on a physical robot requires a laborious setup of the environment and raises significant safety concerns. Therefore, conducting these executions in simulation is a more efficient approach, although the fidelity of the simulator will inevitably cause a sim-to-real gap.
@@
Teleoperation is one of the most common ways to directly capture robot data commanded by human experts. A main advantage of teleoperation is its ability to provide smooth, natural, and precise trajectories for a wide range of tasks. Fig. 11 outlines the control flow of teleoperation data used as a source of policy training. The first step of this process is generating the demonstration through teleoperation, represented by the top path of the control flow (dashed lines). Motion retargeting maps human measurements from the teleoperation device to the desired trajectories in the robot domain. Robot execution data collected from teleoperation can be used to train autonomous policies (solid lines) that directly command the robot’s motion without human intervention.
However, teleoperation for the full-body of a humanoid has a number of limitations. Firstly, a majority of teleoperation systems capture only manipulation skills [10]. Generally, Virtual Reality (VR)-based teleoperation schemes cannot sense the operator’s gait and are restricted to simply commanding walking speeds and directions via joysticks [242, 309, 310]. Full-body sensing, including human gaits, often requires additional equipment such as IMU suits [240, 311] or exoskeletons [245]. These devices may be prohibitively expensive, bulky, complicated to maintain, and lack transparency and user-friendliness. Secondly, while teleoperation can generate versatile training data for a wide range of tasks, the utility of this data may be limited if the robot’s kinematics fail to seamlessly retarget the reference human motion. Additionally, retargeting dynamic tasks such as walking or pushing an object relies heavily on the dynamic model of the human demonstrator [240] and requires meeting high synchronicity [311] and rich sensory feedback.
Although both teleoperation data and policy execution data are recorded from robots, they exhibit distinct characteristics. During teleoperation, human instructors tend to provide diverse demonstrations even for the same task. Therefore, teleoperation data are often multimodal; that is, given a specific task, there exists a distribution of plausible ways to accomplish the task. In contrast, data from executing a single policy are often unimodal; that is, given an input, the output is often fixed. Different policy learning approaches have been proposed to address these multimodal and unimodal data features.
@@ VII-B2 Approaches to Learning from Robot Experience Data
From unimodal policy execution data, which contains paired observations and actions, IL approaches are often used for policy distillation. Behavior Cloning (BC) casts IL as a supervised learning problem, which remains one of the most straightforward approaches for robot skill learning [312, 313]. Another IL technique is Inverse Reinforcement Learning (IRL), which reconstructs rewards from the data in addition to training an RL policy. The IRL study in [300] infers a generalizable reward of the expert demonstration for bipedal locomotion and then uses it to train an RL policy in unseen terrains.
To capture the data multimodality and produce diverse future actions, the Action Chunking Transformer (ACT) [268, 301] is adopted to handle distribution shifts due to the compounding error inherent in naive BC. Diffusion policy [283], a BC method, shows the ability to acquire multimodal locomotion skills by learning from a large dataset collected from multiple expert policies. However, obtaining these skills requires large-scale versatile data, which motivates the scaling of data collection via teleoperation in industrial companies, e.g., Tesla and Toyota Research [10].
Key Takeaways: Although collecting high-quality data demands considerable effort and resources, IL from robot experience remains a reliable method for attaining skills with expert-level performance. Industrial companies and research labs are increasingly focusing on scaling data collection to develop a broader range of diverse policies through IL. Especially, teleoperation is one of the most popular ways to collect humanoid robot experiences nowadays. For further reading on collecting robot experience data, we recommend the survey on humanoid robot teleoperation [314]. We also find the survey on imitation learning of humanoid bipedal locomotion [18] a decent summary.
@@ VII-C Skill Learning: Imitation from Human Data
While robot experiences can serve as a reliable data source, collecting loco-manipulation data directly from robots remains a formidable challenge. Collecting teleoperation data, even though it is one of the most commonly used approaches, is costly and tedious to scale. Additionally, gathering robot data by deploying existing model-based methods or trained policies presents additional difficulties. Deploying these methods on hardware raises safety concerns. On the other hand, data collected from a simulator or synthesized from a model suffers from sim-to-real gaps. Furthermore, model-based methods based on human knowledge (e.g., dynamics models, heuristic trajectories) generate consistent but similar behaviors, leading to limited data diversity.
Learning from a large, diverse corpus of human data mitigates these challenges, as recording human data is more accessible and scalable due to the reflexive usage of loco-manipulation skills by humans in their daily lives. Furthermore, training policies to imitate human data can simplify the synthesis of loco-manipulation behaviors. Recent research efforts in 3D human motion data archival have surged in the vision and computer graphics communities. As shown in Fig. 12, there are currently two primary approaches to acquiring 3D human motion data: (i) recording directly from motion capture systems and (ii) reconstructing from 2D videos.
@@ VII-C1 Obtaining Human Data
Various tracking systems are used to obtain human motion data. As shown in Fig. 13[315, 177] captures humans interacting with various objects while moving around. The following datasets, CMU [316], SFU [317], LAFAAN [318], and AMASS [319], are commonly used because they provide a wide variety of human motions. However, motion capture data often require heavily instrumented environments and actors, making them expensive to scale, and indoor lab settings provide little exposure to outdoor activities.
Alternatively, videos and images obtained from the Internet offer a rich and diverse source of human motion data, including athletic performances, artistic dances, or daily chores. However, motion extracted from internet data is usually of lower quality, containing noise, jittery, and non-physical artifacts due to occlusion and motion blurs. Therefore, the reconstruction of accurate 3D human poses from 2D data remains an active research topic in the computer vision community.
@@
Animation is also a widely explored approach to obtaining human motion data. Although the approach is effective in designing expressive motions for virtual human characters, the process often requires the use of sophisticated animation tools by professional animators [320], which makes it less scalable than motion capture or Internet videos. To address this limitation, researchers have been pursuing motion generation, leveraging human data to generate diverse and realistic motions. Recent advances in generative models, particularly diffusion-based approaches for video generation such as Sora [321] and Stable Video Diffusion [322], have significantly propelled progress in motion generation. A comprehensive survey [323] highlights the impressive capability of these models to generate realistic human motions. However, our focus here is training IL policies to generate physically plausible actions that achieve the motions demonstrated in human data. The training pipeline is shown in Fig. 12.
@@ VII-C2 Challenges in Learning from Human Data
Using offline human motion data (from any source) to train humanoids inevitably creates an embodiment gap both in observation and action spaces due to different body proportions, joint configurations, and mass distributions between most humanoids and humans. Closing this embodiment gap requires retargeting, which involves mapping the motion collected from a source skeletal model to a target robot model. Previous work in computer graphics and robotics has explored various retargeting strategies, such as joint-space [240, 268, 324] and task-space correspondence [301], contact points [219], fingertips [325], gait synchronization [245, 311], and motion feasibility filter [267, 158]. Developing systematic solutions for retargeting the entire human body, including dexterous hands, remains a critical topic for advancing humanoids.
Another correspondence problem is that the human data are only proprioception-based, which lacks sensory input and action output. Notably, these human data lack tactile or force measurement from interactions, which limits the capability of learning for loco-manipulation with rich physical interaction. To solve this issue, IL trains control policies that track references within a physical simulator. Specifically, IL control policies accept state-only references instead of state-action pairs and output control signals. The physical simulator plays a key role in providing sensory input and validating the physical feasibility of the policy action.
@@ VII-C3 Approaches to Learning from Human Data
Examples of learned human-like motions include walking using Generative Adversarial Imitation Learning (GAIL) [307, 303] and Adversarial Motion Prior (AMP) [326, 304, 305], which is an extension of the GAIL. Recently, RL-based motion imitation using motion capture data has successfully transferred to humanoid robots [324, 267, 268, 327]. However, most of these works achieve only relatively conservative loco-manipulation behaviors. Highly agile behaviors, as shown in DeepMimic [328], still exist only in simulators, and similar capabilities have yet to be replicated with real robots.
Human motion imitation can also achieve robust policies capable of rich interactions with objects in unstructured environments. For example, [306, 329] demonstrates the learning of full-body gymnastic skills for humanoids in physics simulation, utilizing video reconstruction data and RL-based motion imitation. [330, 331] learns loco-manipulation policies enabling simulated humanoid characters to carry boxes. Furthermore, [302] mimics motion capture data to achieve tasks such as playing basketball and grasping objects. However, many of these approaches rely on privileged information, such as the ground truth of object and ego poses, from the simulator, which limits their extension to real-world hardware.
Key Takeaways: Transferring skills from humans, especially through Internet-scale datasets, unlocks a broad range of loco-manipulation capabilities for humanoid robots. Although learning from human data holds great promise, a large body of what has been achieved in simulation has yet to be realized in real-world robots. Developing affordable and capable humanoid robots and high-fidelity simulators of real-world loco-manipulation scenarios can accelerate progress in this research direction. As we witness an increasing accessibility of humanoid robots in the market, we foresee that imitation from human data will enable humanoid robots to acquire a vast and diverse skill set. This skill set is fundamental to building humanoid foundation models discussed in Sec. VIII.
@@ VII-D1 Combining Pure RL and IL
The combination of IL and pure RL without demonstration data has led to effective sim-to-real transfers. A two-stage teacher-student paradigm is widely adopted [40, 42, 332]. In these works, a teacher policy is first trained from simulated privileged observation using pure RL. Then, a student policy clones the behavior of the teacher, achieving a similar performance using only partial observations. The trained student policy is readily deployable on hardware with onboard observations. Another two-stage paradigm [333] reverses the order of the two policies. It first uses IL to pre-train an imitation policy from expert data and then uses an RL policy to fine-tune the imitation policy, achieving performance beyond the IL expert and adapting to different environments or tasks [334].
@@ VII-D2 Learning to Track Trajectory Reference
Combining model-based methods with learning-based methods is extensively explored. Numerous works on quadruped robots [335, 336] use MPC to generate reference motions and use them as imitation rewards for RL-based motion imitation. However, calculating MPC online can prolong the training time and occasionally fail to find a feasible solution. Online RL that tracks offline-generated trajectories avoids this problem and has achieved effective RL policies for versatile bipedal locomotion [337, 38, 338]. Supervised learning is also used to mimic TO-generated offline motions. For example, [339] allows humanoid robots to achieve brachiation on monkey bars. In addition to enhancing learning methods with model-informed trajectories, a learned policy can also suggest effective warm starts or hyperparameters for model-based methods, thus significantly reducing the iterations required for convergence to an optimal trajectory.
@@ VII-D3 Learning to Augment Trajectory Reference
Rather than tracking the reference trajectory, augmenting the reference with a residual is another popular approach. Early work on dynamic movement primitives [340] modifies the reference trajectory by learning a task-specific force output, which achieves a humanoid racket swing. A milestone in trajectory augmentation is the demonstration of bipedal locomotion in [341]. Since then, there have been more studies on trajectory augmentation with successful dynamic locomotion [342, 343]. Recently, loco-manipulation skills are also achieved through augmenting TO-based reference trajectory [344]. Beyond joint-level reference, a policy can also augment the task-space reference. For example, [345] learn a locomotion policy to adapt the foot placement reference derived from inverted pendulum models.
Both imitation and augmentation of the reference trajectory share benefits and drawbacks. As a benefit, references expedite the learning process and provide an effective way of learning complex skills. However, both methods rely on predefined trajectories and, therefore, have limited potential to learn emergent, diverse behaviors.
Key Takeaways: This subsection summarizes hybrid methods with successful humanoid hardware deployments. Hybrid methods that combine learning methods with and without demonstration are effective for learning complex and robust behaviors. We also show that model-based methods are not in conflict with learning-based methods, as they can provide guidance for efficient learning. Overall, hybrid methods that combine learning-based and model-based approaches have achieved superior efficiency, versatility, and performance in humanoid tasks that outperform single methods alone.